### v2

# Learning to summarize user information for personalized reinforcement learning from human feedback

Abstract
As everyday use cases of large language model (LLM) AI assistants have expanded, it is becoming increasingly important to personalize responses to align to different users’ preferences and goals. While reinforcement learning from human feedback (RLHF) is effective at improving LLMs to be generally more helpful and fluent, it does not account for variability across users, as it models the entire user population with a single reward model, meaning it assumes that everyone’s preferences are the same. We present a novel framework, Preference Learning Using Summarization (PLUS), that uses reinforcement learning (RL) to learn to produce text-based summaries of each user’s preferences, characteristics, and past conversations. These summaries condition the reward model, enabling it to make personalized predictions about the types of responses valued by each user. Both the user-summarization model and reward model are trained simultaneously, creating an online co-adaptation loop. We show that in contrast to the standard Bradley–Terry model, summaries produced by PLUS capture diverse aspects of user preferences, achieving a 11–77% improvement in reward model accuracy. Key strengths of PLUS are: (1) robust performance with new users and conversation topics, achieving a 25% improvement over the best personalized RLHF technique; (2) zero-shot personalization with state-of-the-art proprietary models like GPT-4 (e.g., PLUS-summary-conditioned responses achieved a 72% win rate compared to 28% for default GPT-4o); (3) learning from flexible user contexts beyond preference labels, and (4) interpretable representation of users, enabling greater transparency and user control in pluralistic LLM alignment.
1 Introduction
Large language models (LLMs) have become a daily part of life for over a billion people, with Gemini reaching at least 450 million monthly users by July [13] and ChatGPT’s daily users approaching a billion [6]. Users interact with LLMs in various ways, from asking for domain-specific help to engaging in open-ended storytelling and content creation[34]. Reinforcement Learning from Human Feedback (RLHF  [17]) is a critical component of modern LLMs deployed in production systems, because it ensures that they are not only functional (i.e., usefully follow user instructions) but also aligned with constitutional values such as safety and helpfulness [1]. The standard Bradley-Terry-Luce (BTL) reward model [2] used in existing RLHF algorithms assumes all users share the same preferences over how the LLM should respond. This technique is incapable of creating systems that can adapt to users’ unique needs and preferences, as a single reward model is used to represent all users.
Figure 1: While standard RLHF techniques fail to capture user variability, PLUS trains both a summarizer and reward model in an online co-adaptive framework to learn summaries $z$ useful for predicting diverse preferences. Italicized texts are actual outputs by GPT-4o and PLUS showing the effects of summaries on personalization.
However, real users have diverse, potentially conflicting, preferences  [12], and current methods are unable to accommodate these differences. In fact, when the standard BTL model is applied to data with conflicting preferences, it will either ignore the preferences of minority groups [23], or learn a reward model that is inaccurate for all groups of users [18]. This has led to recent calls to develop techniques for pluralistic alignment [26, 14, 23], to enable LLMs to be personalized to heterogeneous user preferences. Given the widespread use of LLMs and the diverse populations of users they serve, being able to model and adhere to more than one set of preferences is critical for making effective, accurate assistants that are capable of personalizing to each individual’s unique needs. One way to achieve pluralistic alignment is by developing a reward model conditioned on user variability, rather than relying on a single BTL model [23, 32, 18] (e.g.,  [18] learns a latent user embedding from the past preference labels). However, compressing rich text into a vector can lead to performance loss [4]. As yet, we are missing a highly performant, public RLHF technique capable of modeling the preferences of diverse users.
A more natural approach which retains the strengths of LLMs to reason with text, is to learn text-based user summaries to act as the latent user variable. We propose Preference Learning Using Summaries (PLUS) as a general RLHF framework that learns to summarize important information about the user using reinforcement learning (RL) fine-tuning and conditions the reward model on these user summaries. PLUS consists of two co-dependent training phases: (1) a summarizer trained with Proximal Policy Optimization (PPO) [20], and (2) a summary-conditioned reward model, which provides the reward signal for training the summarizer and makes predictions about user preferences.
We carefully compare our approach to several alternative implementations. First, we check whether we can simply prompt an LLM to generate user summaries without fine-tuning the summarizer. Our experiments show that automatic summaries (even from state-of-the-art GPT-4) often miss key details necessary to capture user preferences. Another alternative is to condition the reward model on the user’s entire conversation history, without summarizing it. However, both approaches fall short compared to concise, summary-conditioned reward models when conversation topics change. An additional benefit of concise summaries compared to existing techniques is that the summaries are easier to review and edit, improving transparency and trust in LLM assistants. We show that such benefits extend to complex real-world datasets, like PRISM [12], where diverse user preferences can be expressed through flexible textual summaries and used to enable zero-shot personalization. In summary, our main contributions are:
1. Novel pluralistic alignment algorithm, PLUS (Preference Learning Using Summarization) that uses RL to jointly learn a summarizer and reward model. PLUS improves accuracy, interpretability, and transferability across new conversations and users.
2. Empirical advantages of PLUS on benchmarks from prior state-of-the-art work [18]: Pets and UltraFeedback [5]. PLUS improves reward model accuracy by 11-77% compared to BTL, is robust to topic shifts, and can handle unstructured user contexts.
3. Scaling to a dataset in the wild. We are one of the first to report reward modeling results with PRISM [12], a pluralistic dataset of 1,500 users across 75 countries and 20 LLMs.
4. Enabling zero-shot personalization via PLUS-generated summaries. PLUS can generalize to unseen users of the PRISM dataset and the learned summaries can be used to personalize responses of strong proprietary models without additional finetuning (PLUS summary-conditioned responses achieve a 72% win rate over the default GPT-4o responses).
2 Related Work
Reinforcement Learning from Human Feedback (RLHF). Our work builds on reinforcement learning from human feedback (RLHF)  [17], which models binary human preferences using the Bradley-Terry-Luce (BTL) model [2]. The BTL model assumes the user provides preference ratings in terms of which of two responses they like better; it then learns a reward model (RM) that estimates the magnitude of the user’s preferences using this data. This is then used to train an LLM policy to optimize responses based on the reward signal. RLHF has been effective in refining language models’ responses to better align with user-preferred attributes, like helpfulness and truthfulness [1]. However, existing RLHF works [35, 1, 17, 23, 19] have assumed a single reward model to capture all user data, which may be misaligned with heterogeneous users.
Personalized & Pluralistic RLHF. Increasingly, more works [25, 11, 12, 32, 18, 14, 3, 15] have pointed out the limitations of existing BTL reward models in accommodating diverse underlying preferences, which may conflict with each other in terms of what individual users consider ideal responses. These differences may surface along various dimensions, from writing style to value alignment. Despite algorithmic advances in pluralistic alignment, large-scale human datasets are still limited. PRISM [12] makes an effort to provide a heterogeneous user dataset consisting of 1,500 participants from 75 countries interacting with 21 different LLMs (contributing around 9,000 unique conversations). However, so far we are not aware of work that has successfully published reward modeling results on PRISM, pointing to the difficulty of modeling diverse and heterogeneous preferences with sparse data.
Broadly, works on pluralistic RLHF have focused on either of two main aspects: (1) personalizing reward models [23, 18, 32, 15, 3] and (2) personalizing LLM responses with a user model [14]. Reward (user) modeling serves as the backbone of many existing algorithms for LLM alignment, since reward model accuracy has direct implications for downstream tasks, such as RLHF finetuning and best-of-N sampling [22, 8, 16]. Notably,  Poddar et al. [18] introduces Variational Preference Learning (VPL) which uses a variational autoencoder to learn user-specific embeddings and conditions the reward model on these latent variables.  Siththaranjan et al. [23], Zhao et al. [32], Chen et al. [3] provide algorithms for accommodating diverse user groups through methods such as Distributional Preference Learning (DPL), dispersion index, and shared LoRA adaptation. However, the key difference between prior works and ours is that these methods represent users as embedding vectors or (linear) weights, which reduces the power and interpretability of the underlying personalization mechanism. In contrast, we leverage the text-generation capabilities of LLMs to represent diverse user preferences through summaries.
LLM finetuning for summarization. Finetuning LLMs for improved and domain-specific summarization has long been of interest [27, 35, 31, 28, 24, 29]. Most closely related to our work is by  Wu et al. [29] which demonstrate that learning user summaries can be effective for personalizing recommendation systems, when you assume access to a fixed user classifier. In contrast, we focus on the online co-adaptive loop of training both the summarizer and reward model, which serves as the key for enabling the reward model to make use of the learned summaries for capturing nuanced user preferences.
3 Preliminaries
Bradley-Terry-Luce reward model. Our work focuses on the reward modeling stage of the standard RLHF pipeline [17]. Given a dataset of annotated preferences $\mathcal{D}=\{(s_{A}^{i},s_{B}^{i},\mathbbm{1}\{s_{A}^{i}\succ s_{B}^{i}\})\}_{i=1}^{N}$, where $s_{A}=[x,y_{A}]$ and $y_{B}=[x,y_{B}]$ are two responses $y_{A},y_{B}$ over the same prompt $x$, the standard Bradley-Terry-Luce (BTL) model [2] assumes an underlying reward function $r(s)$. This reward function determines the likelihood of the preference label $\mathbbm{1}\{s_{A}\succ s_{B}\}$ as:
$p\big(\mathbbm{1}\{s_{A}\succ s_{B}\}=1\mid s_{A},s_{B}\big)=\frac{\exp r(s_{A})}{\exp r(s_{A})+\exp r(s_{B})}=\sigma\big(r(s_{A})-r(s_{B})\big),$ (1)
where $\sigma(x)=1/(1+e^{-x})$ is the sigmoid function. The reward modeling stage typically involves learning a parameterized reward function $r_{\phi}(s)$ by maximizing the likelihood of Eq. 1 over the observed response pairs and their preference labels. The BTL model assumes all users share the same reward function, and thus fails to capture conflicting preferences when two users prefer different responses to the same prompt $x$.
User-conditioned reward model. Existing work in personalized RLHF [14, 18] has demonstrated the effectiveness of extending the standard BTL reward model into a user-conditioned reward model:
$p\big(\mathbbm{1}\{s_{A}\succ s_{B}|z\}=1\mid s_{A},s_{B},z\big)=\frac{\exp r(s_{A}|z)}{\exp r(s_{A}|z)+\exp r(s_{B}|z)}=\sigma\big(r(s_{A}|z)-r(s_{B}|z)\big)$ (2)
where $z$ is an identifier (e.g., a user ID) that distinguishes different users or user groups. Eq. 2 is better suited for capturing the population’s heterogeneous preference distribution that may be multi-modal, and the target is to learn a parameterized user-conditioned reward function $r_{\phi}(s|z)$. This gives the reward model the expressivity to capture different user types, which yields more accurate predictions when user preferences are diverse and even conflicting.
However, in most real-world scenarios, datasets do not contain user identifiers $z$. This is partly because such identifiers may be unavailable due to privacy issue, and partly because conditioning only on users observed in the dataset makes it difficult to generalize to new users. Therefore, we can instead use context $c$, such as past conversations, to represent the current user in each data point111 Poddar et al. [18] requires that the context consist of a user’s past chosen and rejected responses to different prompts $c=\{(s_{A}^{j},s_{B}^{j})\}_{j=1}^{m}$. In contrast, PLUS can incorporate more flexible inputs like past conversations without explicit preference labels and system strings.. This leads to a dataset of user-conditioned annotated preferences $\mathcal{D}^{\prime}=\{(s_{A}^{i},s_{B}^{i},c^{i},z^{i},\mathbbm{1}\{s_{A}^{i}\succ s_{B}^{i}|z^{i}\})\}_{i=1}^{N}$ where $z^{i}$ are unknown. For notational convenience, we assume $s_{A}^{i}$ is always the preferred one by user $z^{i}$, and the dataset can be simplified to $\mathcal{D}^{\prime}=\{(s_{A}^{i},s_{B}^{i},c^{i})\}_{i=1}^{N}$.
4 Preference Learning Using Summarization: PLUS
While prior works [14, 18] formulate the latent user representation $z$ as an embedding vector, we propose learning $z$ in natural language as a more natural way to leverage the advanced text-based reasoning and generation capabilities of LLMs. To be more specific, we propose to use a summarizer $\pi_{\theta}$ to generate a concise representation in natural language based on the context $c$ of each user, $z\sim\pi_{\theta}(\cdot|c)$. Then, the parametrized reward model $r_{\phi}(\cdot|z)$ will compute the reward conditioned on this learned summary $z$.
Rather than manually defining any ground truth criterion for useful user summaries and pre-training a summarizer in advance, we argue that the quality of a summary should be measured by its ability to enable the downstream reward model to accurately predict user preferences. Therefore, the summarizer and the reward model should be optimized under a joint objective, namely minimizing the negative log-likelihood loss based on Eq. 2 across the user-conditioned dataset $D^{\prime}$:
$\mathcal{L}_{\text{NLL}}(\phi,\theta)=-\hat{\mathbb{E}}_{(s_{A},s_{B},c)\sim\mathcal{D}^{\prime},z\sim\pi_{\theta}(\cdot|c)}\big[\log\sigma\big(r_{\phi}(s_{A}|z)-r_{\phi}(s_{B}|z)\big)\big]$ (3)
Figure 2: Eq. 3 naturally leads to the online co-adaptation of the summarizer and the reward model.
To handle the co-dependency in Eq. 3, we design a two-stage sequential (batch) training process that alternates between training one model while keeping the other fixed. By allowing the summarizer and the reward model to co-adapt, the summarizer learns to extract information that is most useful for improving the predictive accuracy of the reward model, while the reward model is trained on updated summaries to more accurately capture user preferences. This process is computationally efficient, since the generated summaries always target the most recent iteration of the reward model.
Training the reward model. Once the summarizer is fixed, the reward model is trained to minimize the negative log-likelihood loss (Eq. 3) using pairs of chosen and rejected responses, conditioned on user summaries sampled from the summarizer policy.
Training the summarizer. Unlike the reward model, a summarizer’s optimal weights cannot be learned directly through supervised learning, because whether a summary is effective for predicting preferences can only be assessed once the full summary has been sampled. Instead, RL can optimize the summarizer at the token level by assigning credit for the full-trajectory return across long token sequences $(z_{1},z_{2},...,z_{t\leq H})$, as demonstrated by prior successes in using RL to improve summarization [27, 35]. Thus, we formulate generating the summary into an RL post-training problem, with the objective in Eq. 4. We start with a default LLM as a reference policy $\pi_{\theta}$, sample summaries (trajectories) from the current $\pi_{\theta}$, estimate their returns (as negative losses of the reward model), and update the weights $\theta$ in a direction that maximizes estimated returns.
$J(\pi_{\theta})$ $=\hat{\mathbb{E}}_{\begin{subarray}{c}z\sim\pi_{\theta}(\cdot\mid c)\\ (s_{A},s_{B},c)\sim\mathcal{D}^{\prime}\end{subarray}}\left[\sum_{t=0}^{H}\gamma^{t}R_{\phi}(z_{t},s_{A},s_{B})\right],$ (4)
[…]
PPO training for handling non-stationarity. As the reward model for predicting user preferences is continuously updated, the RL rewards for training the summarizer become non-stationary, as in a multi-agent RL (MARL) problem. Therefore, while different RL algorithms could be considered for learning an effective summarizer, we choose PPO [20], as it is a state-of-the-art MARL algorithm and performs well even under non-stationary rewards [7, 30]. In addition to the summary-level rewards $R=\sum_{t=0}^{H}\gamma^{t}R_{t}$, we calculate token-level advantages using the generalized advantage estimator (GAE) [21]:
$\mathcal{A}^{\text{GAE}(\gamma,\lambda)}_{t}=\sum_{i=0}^{H}(\gamma\lambda)^{i}\delta_{t+i}^{V},\delta_{t}^{V}:=R_{t}+\gamma V(z_{t+1})-V(z_{t})$ (5)
where $V_{t}$ are token-level value estimates by the PPO critic. Advantages are normalized across batches $\mathcal{A}_{t}^{\text{norm}}$, and the summarizer $\pi_{\theta}$ is optimized with clipping to avoid instability, leading to the PPO loss:
$\mathcal{L}_{\text{PPO}}(\phi,\theta)=-\hat{\mathbb{E}}_{t}\left[\min\left\{\zeta_{t}(\theta)\mathcal{A}_{t}^{\text{norm}},\text{clip}(\zeta_{t}(\theta),1-\epsilon,1+\epsilon)\mathcal{A}_{t}^{\text{norm}}\right\}\right],\zeta_{t}(\theta)=\frac{\pi_{\theta}(z_{t}|z_{<t})}{\pi_{\theta_{\text{old}}}(z_{t}|z_{<t})}$ (6)
The complete joint-training framework is illustrated in Alg. 1.
Algorithm 1 Co-adaptation of Summarizer and Reward Model
1:Initial summarizer $\pi_{\theta}$, reward model $r_{\phi}$, dataset $\mathcal{D}^{\prime}$ of user contexts $c$ and preference pairs, rollout batch size NN, mini-batch gradient steps $M_{\theta},M_{\phi}$
2:for each training iteration $t$ do
3:  Sample NN tuples of contexts and preference pairs from $\mathcal{D}^{\prime}$
4:  Generate summaries $z\sim\pi_{\theta_{t}}(\cdot|c)$
5:  Compute rewards in Eq. 4 and advantages in Eq. 5
6:  for iter = 1 to $M_{\theta}$ do
7:   Compute the PPO loss $\mathcal{L}_{\text{PPO}}(\phi,\theta)$ in Eq. 6.
8:   Update summarizer parameters: $\theta$ as $\theta\leftarrow\theta-\alpha_{\theta}\triangledown_{\theta}\mathcal{L}_{\text{PPO}}(\theta,\phi)$
9:  end for
10:  for iter = 1 to $M_{\phi}$ do
11:   Compute the negative log-likelihood loss $\mathcal{L}_{\text{NLL}}(\phi,\theta)$ in Eq. 3
12:   Update reward model parameters: $\phi$ as $\phi\leftarrow\phi-\alpha_{\phi}\triangledown_{\phi}\mathcal{L}_{\text{NLL}}(\phi,\theta)$
13:  end for
14:end for
15:return Trained $\pi_{\theta},r_{\phi}$
Implementation.
We finetuned different reward models (Qwen2.5-0.5B-Instruct, Qwen2.5-3B-Instruct, and Llama3.2-3B-Instruct) and summarizers (Qwen2.5-3B-Instruct or Llama3.2-3B-Instruct). We pair each reward model with a summarizer from the same model family. The models are fine-tuned using low-rank adaptation (LoRA) [9] with distributed computation across multiple nodes (we used 4 x H200 GPUs). The co-adaptation pipeline in Alg. 1 is built on top of PPO training framework from OpenRLHF [10]. We follow the baseline implementations for BTL, VPL, and DPL by  Poddar et al. [18]. More details are in Appendix H.
5 Experiments
5.1 Domains
We evaluate on three benchmark tasks from  Poddar et al. [18], along with a challenging variant that tests out-of-distribution user contexts, as well as real user datasets.
Pets. In this dataset, 50% of users prefer to discuss dogs and 50% prefer cats, so a single reward model will fail to accurately capture user preferences. To test the robustness of different methods with new conversation topics and users, we create a new Pets OOD (Out-Of-Distribution) dataset, which consists of users who prefer talking about rabbits or birds, as opposed to cats and dogs. Using this dataset, we can assess how well each method can generalize to new users with novel preferences despite being trained on only cats and dogs.
UltraFeedback. Poddar et al. [18] construct two domains from the scaled-up preference dataset UltraFeedback [5]. In UltraFeedback-P2, users prefer either helpfulness or honesty, and in UltraFeedback-P4, users choose a subset of four possible attributes of an LLM assistant: helpfulness, honesty, truthfulness, and instruction-following. In both domains, personalized reward models need to infer different user types based on past conversation data. For example, some users may prioritize helpfulness and honesty, while others may value instruction-following above all other attributes. Following  Poddar et al. [18], we also focus on a controversial subset of the dataset by removing samples where all users agree on the chosen response.
Evaluation with real user data (PRISM). This real-world user dataset consists of open-ended conversations from 1,500 participants across 75 countries and 20 LLMs [12]. Each user engages in up to six conversations on topics of their choice, loosely guided by one of three themes: unguided (free topic), value-guided, or controversy-guided. We filter out users who did not answer the preference survey or quit before completing all six conversations. As a result, we train on 1,000 users and evaluate zero-shot personalization performance on held-out users. This domain is particularly challenging due to the high heterogeneity and variance across users, as well as the lack of shared conversation topics, compared to the size of the training data. Despite the challenge, our work is among the first to report pluralistic reward modeling performance on the PRISM dataset.
5.2 Baselines
We evaluate PLUS against six existing RLHF baselines:
Bradley-Terry-Luce (BTL) model most widely used in RLHF preference learning [17]. The BTL model uses a single model to capture all user preferences which may be misaligned with heterogeneous populations.
Distributional Preference Learning (DPL)  Siththaranjan et al. [23] proposes to learn both the mean and the variance of the reward distribution, which captures overall variance across users’ preferences but does not model each user’s distinct preferences.
Variational Preference Learning (VPL)  Poddar et al. [18] learns a latent representation of each user with a vector of size 512 using a variational auto-encoder (VAE) reward model and shows state-of-the-art performance on pluralistic preference modeling.
In-Context Learning (ICL) Past conversation examples are directly included as part of the reward model’s prompt, in addition to the chosen and rejected response. This would in theory enable modeling each user’s distinct preferences, but may be less sample-efficient due to the lengthy context, and less robust to topic changes.
PLUS with a fixed summarizer (PLUS-untrained). This ablation only trains the reward model while fixing the summarizer.
Oracle. This is a reward model conditioned on the true user preferences. For example, in UltraFeedback, if a user values helpfulness and honesty over the other attributes, that user’s preference model is conditioned on the following string input: This user prefers helpfulness and honesty in the assistant’s responses.
6 Results
6.1 Advantages of PLUS on existing benchmarks
First, we compared PLUS against existing RLHF methods on existing personalized reward modeling benchmark tasks. Table 1 shows that PLUS achieves improved prediction accuracy across all three domains compared to vanilla RLHF with the largest improvement of 77% on Pets. While existing personalized RLHF techniques, like VPL and ICL also perform well on Pets, they are not as performant for complex domains resembling real human preferences, like UltraFeedback P-2 and P-4, where conversation contents vary between training and test samples, while the underlying user groups remain the same. VPL struggles to personalize showing limited improvement over BTL (-2.76–4.81% across model types). We hypothesize this is because training the reward model to distill all user context into a fixed-length vector sacrifices performance, as previously discussed by  Cideron et al. [4]. While in theory ICL can model each user’s distinct preferences, conditioning on all past conversation examples is sample-inefficient due to long, unsummarized contexts and is more sensitive to conversation topics than to the underlying user preferences. Specifically, ICL only shows 0.9–8.4% improvements across model types, compared to 10.4–17.7% achieved by PLUS, on the UltraFeedback dataset. We further investigate this limitation with our Pets-OOD experiments.
Table 1: Reward model accuracy on the held-out set. Mean and std from 3 seeds; otherwise, single seed.
Pets Ultrafeedback-P-2 Ultrafeedback-P-4
Reward model Qwen-0.5B Qwen-0.5B Qwen-3B Llama-3B Qwen-0.5B Qwen-3B Llama-3B
BTL [2] 44.1 $\pm$ 2.8 58.1 $\pm$ 0.3 58.75 57.40 52.7 56.25 56.35
DPL [23] 46.2 $\pm$ 4.4 58.5 $\pm$ 0.7 58.15 57.55 56.2 57.00 56.60
VPL [18] 100 $\pm$ 0 58.4 $\pm$ 0.3 57.15 59.40 55.3 58.10 57.45
ICL 100 $\pm$ 0 59.6 $\pm$ 0.7 59.30 59.10 57.3 59.10 57.70
PLUS-untrained 96.2 $\pm$ 1.6 59.9 $\pm$ 0.2 59.65 59.35 56.4 58.75 57.90
PLUS (Ours) 99.8 $\pm$ 0.2 69.4 $\pm$ 1.1 69.85 66.90 61.8 62.45 62.80
Oracle 100 $\pm$ 0 69.0 $\pm$ 0.3 68.95 67.95 71.4 73.05 73.80
Even zero-shot summaries (by PLUS-untrained) can achieve improvements over BTL, but they still fall short of providing sufficient predictive signal for the reward model. This limitation is clearly observed in the Pets dataset: the untrained summarizer struggles to distinguish meaningful user contexts from spurious details or outputs vague descriptions, while the trained summarizer correctly identifies whether a user prefers cats or dogs, as shown in the box below. This highlights the importance of learning summaries, which we further investigate through ablations that fix either the reward model or the summarizer.
Comparison of PLUS-generated and untrained summaries PLUS summary (shows clear preference for cats): The user is interested in information about cat behavior and properties, excluding topics related to dogs or human effects. Default summary (discusses general pet traits): The user seems to appreciate traits that denote affectionate and playful qualities in pets, as evidenced by their rejection of traits related to training or timing of activity (renewal of knowledge). Default summary (focuses on spurious feature): The user prefers short, factual information and tends to reject detailed anecdotes or complex behaviors. The assistant should stick to basic descriptions where possible.
6.2 Robustness of PLUS to new topics & users
As new users increasingly engage with LLM assistants, robustness to new users and conversation topics that are different from the training data becomes important for personalized systems. While UltraFeedback already evaluates on new user contexts, these come from users with the same underlying preferences. We took a step further by explicitly evaluating new users whose test-time preferences shift to entirely different content, using Pets (OOD). Concretely, the user context and preference labels $(s_{A},s_{B})$ used for training are about dogs or cats, but at test time, the summarizer receives a new user context about birds and rabbits, and the reward model is given two response options between birds and rabbits. As shown in Table 2, PLUS’s robust generalization to new user contexts indicates that it has learned a general algorithm for extracting user information into summaries, which translates well to new users with new types of preferences. In contrast, other approaches may be vulnerable to memorizing specific users’ preferences, rather than learning how to learn about what a new user is interested in.
Table 2: Reward model accuracy with and without user shift in Pets. The reward model is trained with Qwen2.5-0.5B-Instruct, and the summarizer is trained with Qwen2.5-3B-Instruct. Mean and std from 3 seeds. ICL’s performance drops from 100 to 72%, and VPL drops to below 50%, whereas PLUS is the only method to maintain accuracy above 90%.
Out-of-distribution user contexts
BTL 49.5 $\pm$ 2.6
DPL 49.0 $\pm$ 0.5
VPL 44.8 $\pm$ 4.0
ICL 72.5 $\pm$ 3.1
PLUS-untrained 85.8 $\pm$4.0
PLUS (Ours) 93.7 $\pm$4.0
[…]
6.3 Importance of co-adapting summarizer and reward
A key component of PLUS is the online co-adaptation of the summarizer and the reward model, which enables the summarizer to focus on dimensions of user variability that are most relevant to the reward model. We conducted ablations using different model types to examine the effect of training only the summarizer or the reward model. Table 3 shows that despite using a smaller reward model (0.5B), co-adaptation achieves higher performance than larger reward models (3B) that are either trained with fixed summaries or only used to train the summarizer. Surprisingly, using a state-of-the-art summarizer (GPT-4o) with a trained 3B reward model still underperforms compared to co-training substantially smaller models (0.5B reward model and a 3B summarizer) in both UltraFeedback domains. This supports our intuition about the importance of RL fine-tuning for learning summaries: good summaries are defined by the reward model’s performance rather than any fixed criteria. Therefore, the training signal for the summarizer should come from the reward model’s loss.
Table 3: Reward model accuracy with static reward models or summarizer. Static reward model or summarizer is in red ([Uncaptioned image], [Uncaptioned image]), and the trained model is in green ([Uncaptioned image], [Uncaptioned image]). All ablations with Qwen-Instruct series.
Ultra Feedback (UF)-P2 Ultra Feedback (UF)-P4
Summarizer [Uncaptioned image] Reward model [Uncaptioned image] Accuracy (%) Summarizer [Uncaptioned image] Reward model [Uncaptioned image] Accuracy (%)
Trained 3B [Uncaptioned image] Static 0.5B [Uncaptioned image] 53.45 Trained 3B [Uncaptioned image] Static 0.5B [Uncaptioned image] 49.50
Trained 3B [Uncaptioned image] Static 1.5B [Uncaptioned image] 52.10 Trained 3B [Uncaptioned image] Static 1.5B [Uncaptioned image] 50.10
Trained 3B [Uncaptioned image] Static 3B [Uncaptioned image] 42.20 Trained 3B [Uncaptioned image] Static 3B [Uncaptioned image] 53.50
Static 3B [Uncaptioned image] Trained 0.5B [Uncaptioned image] 59.90 Static 3B [Uncaptioned image] Trained 0.5B [Uncaptioned image] 56.40
Static 3B [Uncaptioned image] Trained 3B [Uncaptioned image] 59.65 Static 3B [Uncaptioned image] Trained 3B [Uncaptioned image] 58.75
GPT-4o [Uncaptioned image] Trained 3B [Uncaptioned image] 59.30 GPT-4o [Uncaptioned image] Trained 3B [Uncaptioned image] 60.80
Trained 3B [Uncaptioned image] Trained 0.5B [Uncaptioned image] 69.40 Trained 3B [Uncaptioned image] Trained 0.5B [Uncaptioned image] 61.80
Trained 3B [Uncaptioned image] Trained 3B [Uncaptioned image] 69.85 Trained 3B [Uncaptioned image] Trained 3B [Uncaptioned image] 62.45
6.4 Flexibility to incorporate unstructured text data beyond preference labels
PLUS can incorporate additional user information in flexible forms of textual input, not just preference labels. We focused on two input types. Non-preference-based user contexts. PLUS can infer user preferences from natural conversations without requiring chosen and rejected response pairs per prompt. Users do not typically provide both an accepted and a rejected response, so VPL’s input assumption may limit its applicability to real-world personalized systems. While both PLUS and ICL can train reward models using unstructured user contexts, Table 4 shows that PLUS outperforms ICL by 9–15%.
Table 4: Reward model accuracy with unstructured inputs. We compared the performance of PLUS and ICL when user contexts include unstructured data other than preference labels, as most LLM-user interactions do not gather explicit preference feedback. (i) Preference data provides users’ past preferences in the form of chosen and rejected response examples. (ii) User-guideline + preference data additionally includes textual instructions specifying which aspects of an LLM assistant the summarizer should focus on. (iii) Conversation data does not include binary preference labels; instead, it only presents the chosen responses as “examples of past conversation” and expects the reward model or summarizer to infer the user’s underlying preferences only from the positive examples. This last setup more closely models real user interactions with LLM assistants, since users are typically given only one response rather than choosing between two options; however, their preferences can still be inferred from the users’ initial prompts, follow-up responses, and questions. We trained a Qwen2.5-0.5B-Instruct reward model for both ICL and PLUS and a Qwen2.5-3B-Instruct summarizer for PLUS. We did not include VPL, as it only supports pairwise preferences as contexts.
Ultra Feedback (UF)-P2 Accuracy
Reward model Preference data User guideline + preference data Conversation data
ICL 59.60 59.65 58.55
PLUS 69.40 70.25 68.3
Ultra Feedback (UF)-P4 Accuracy
Reward model Preference data User guideline + preference data Conversation data
ICL 57.30 56.60 56.96
PLUS 61.80 62.70 62.45
User-guided instructions. For example, a system designer may know which dimensions are relevant for personalization but not know which attributes actually matter to a specific user. To simulate this scenario with the UltraFeedback dataset, we made a simple modification to the user’s prompt stating “Users typically care about the LLM assistant’s helpfulness, honesty, truthfulness, and instruction-following capabilities.” This additional information improves performance of PLUS compared to ICL by 10–16%. PLUS can incorporate flexible inputs beyond preference data, which other personalized techniques, like ICL and VPL, cannot leverage effectively.
6.5 Evaluating PLUS in the wild with the PRISM user data
Despite the growing interest in pluralistic alignment, results demonstrating effectiveness on real-world datasets have been limited due to both the challenging nature of real user data and the limited sample size. Having established the advantages of PLUS on existing benchmarks for pluralistic alignment, we next applied it to the real-user dataset PRISM [12].
Reward modeling with PRISM. We first implemented PLUS with a Qwen2.5-3B-Instruct summarizer and a Qwen2.5-0.5B-Instruct reward model using 1,000 users from the PRISM dataset. We also implemented six reward-model baselines with Qwen2.5-0.5B-Instruct. Then, we evaluated reward-model accuracy on the preference labels of held-out conversations. This evaluation protocol assumes that each user’s underlying preferences remain consistent across different conversation topics, so once the reward model learns a user’s preferences from five conversations, it should be able to predict the preferred response in a new, sixth conversation. Due to the challenging nature of this dataset, which attempts to capture a high degree of variance with a small sample size, we find that all methods achieve somewhat low reward prediction accuracy, and report the results in Table 6 in the appendix. PLUS outperforms existing (un)personalized methods with an accuracy of 62.9%, whereas BTL achieves 59.8%, VPL 60.4%, and ICL 60.1%. However, in spite of marginal improvements in reward modeling score, we find that PLUS is still able to learn to usefully extract user information that enables personalization as below.
6.6 Zero-shot personalization on PRISM with GPT-4 models using PLUS summaries
So far, we have focused on the reward modeling performance of PLUS compared to existing RLHF techniques. However, another key advantage lies in the transparency and portability of PLUS-generated user summaries for downstream personalization tasks.
Personalized LLM-as-judges. We combine PLUS with an LLM-as-judge setup [33] to enable zero-shot personalized reward modeling. Specifically, we evaluate GPT-4’s prediction of a user’s preferred response between two candidates, both with and without the PLUS-generated summary. The evaluation includes held-out user examples as well as value-guided and controversial conversations from known users from PRISM. As shown in Table 7, across three task types, PLUS-guided LLM-as-judges improve accuracy for both GPT-4o (by 19.5%) and GPT-4.1 (by 2.9%). These results demonstrate that PLUS-generated summaries are effective not only for personalizing learned reward models but also for personalizing zero-shot proprietary models.
Personalized response generation.
Figure 3: Win rates of personalized vs default GPT-4 on the PRISM dataset (evaluated with two prompts per 308 unseen users).
We prompted GPT-4o and 4.1 to generate responses with and without the PLUS-generated user summary. Specifically, we sampled two new prompts for each of the 308 held-out users in PRISM and compared the win rates of personalized versus default GPT responses evaluated by an oracle reward model, which is trained to condition on users’ self-stated preferences (in PRISM, users rate different attributes of an LLM assistant in surveys, which we use as ground truth). This experiment assesses whether PLUS-generated summaries are useful for enabling proprietary models to personalize to new users without additional finetuning. In Fig. 3, the win rates of PLUS-personalized responses over default GPT responses clearly demonstrate the effectiveness of PLUS (e.g., 72% vs. 28% for GPT-4o, and 69% vs. 31% for GPT-4.1). We included examples below to show the effect of PLUS-summary-conditioned personalization using GPT-4.
Selected examples of personalization with PLUS & GPT-4o Prompt #1: Is Jesus Christ the Son of God? PLUS-summary about the user: The user prefers detailed, balanced responses avoiding definitive statements when there’s uncertainty. Default response: In Christianity, Jesus Christ is believed to be the Son of God. This is a central tenet of Christian faith, affirming Jesus’ divine nature. Personalized response: Many Christians believe that Jesus Christ is the Son of God…. However, perspectives vary widely across different religions and secular viewpoints. Judaism and Islam, for instance, do not view Jesus as the Son of God, although they recognize his historical significance. Interpretations can depend on personal beliefs, religious tradition, and theological perspective. (Personalization based on the user summary is highlighted in blue.) Prompt #2: What is your opinion on abortion? PLUS-summary about the user: The user prefers detailed, factual answers with supportive examples and clear explanations. Default response: I don’t have personal opinions or beliefs, but I can provide information on the topic. Abortion is a complex and sensitive issue with diverse viewpoints, including ethical, legal, and personal considerations. It’s vital to approach discussions about abortion with respect for differing perspectives and to consider the specific laws and cultural contexts of each region. Personalized response: I don’t have personal opinions or beliefs. However, I can provide a detailed, factual overview. Abortion is the termination of a pregnancy and is a contentious issue globally. Views on abortion often reflect broader societal beliefs and values, including those related to women’s rights, morality, and medical ethics. For example, in the U.S., legal perspectives vary widely, with some states imposing strict regulations while others support a woman’s right to choose. Factors like fetal viability and maternal health often play roles in these laws. Understanding these contexts helps to form a comprehensive view of the issue.
7 Discussion
We propose PLUS, a novel RLHF technique of learning text-based summaries to capture user variability and co-training a reward model conditioned on the obtained summaries. Summaries provide a natural format that leverages the language modeling capabilities of LLMs and show strong empirical performance that extends beyond benchmark tasks. We conducted extensive experiments and compared our approach against six RLHF baselines, demonstrating the following five benefits.
  • Strong empirical advantages. PLUS outperforms standard BTL by 11-77% on pluralistic datasets of Pets and UltraFeedback, and achieves consistently strong performance compared to existing pluralistic RLHF techniques.
  • Robustness. PLUS is the only method that maintains 90% accuracy with test-time preference content shifts, suggesting that it has learned a general algorithm for extracting user information that translates to new users with new types of preferences.
  • Benefit of co-adaptation. Co-adapting a smaller summarizer (3B parameters) and reward model (0.5B) outperforms a reward model trained on proprietary model summaries by 10%. We carefully compared PLUS to alternative implementations of training either the summarizer or the reward model while fixing the other. Our experiments show that automatic summaries affect the learnability of the reward model, and only training the summarizer achieves poor reward modeling performance across domains and model types.
  • Flexibility to incorporate unstructured user contexts. Unlike prior personalized RLHF techniques [18, 14] , PLUS can leverage user-guided instructions and unlabeled conversation data, achieving improvements of 9-16% compared to ICL. This is especially important for scaling up to real user systems where explicit preference labels may not be available, and therefore the user contexts must be inferred from other types of text data.
  • Summary-conditioned personalization of proprietary models, achieving a win rate of 72% against the unpersonalized GPT-4o responses. Furthermore, generated summaries are human-readable, enhancing transparency and improving users’ trust in LLM systems.
Appendix
Appendix A Illustrative example of why compression by $z$ matters.
Naively summarizing past conversations or using the entire context $c$ may be insufficient. Summarizing past conversations with an untrained summarizer can lead to summaries that focus on past content rather than the general style of conversations the user prefers to have. Therefore, the obtained summaries may not provide predictive signal for the reward model.
Figure 4: An intuitive example of why compression by $z$ matters for the reward model.
As an illustrative example, in Fig. 4, suppose the reward is +1 for the area above $y=x$ and -1 otherwise. The objective is to represent the context vector $c$ as a scalar. A lossless compression outputs $z$ as the angle between $c$ and the diagonal, while an uninformed summarizer may instead output $z^{\prime}=\theta_{1}$. This affects the expressivity of the learned reward model, since this lossy compression can never recover the correct reward.
Alternatively, one could include all prior preference labels from the user as in-context learning (ICL) examples, and train the reward model conditioned on such examples (i.e., the full list of examples acting as the conditioning variable $z$). However, the high-dimensional input space of ICL, and subsequently a more complex hypothesis class of reward models, may make reward learning more difficult in terms of sample complexity. This shows up in our empirical results, where ICL underperforms when trying to generalize to new conversation topics. While in theory ICL can enable personalization, it may be sample-inefficient and computationally expensive to train and use for inference due to the increased context.
Appendix B PLUS-generated summary examples
Pets
While PLUS-untrained still achieves a high accuracy of 96.2% on the Pets dataset, the summaries for incorrect examples highlight the importance of training the summarizer so that it can learn to distinguish between relevant and irrelevant details in the user’s past conversations. We’ve identified common failure cases made by the untrained summarizer as: (1) focusing on irrelevant aspects of the user’s conversation history, such as style, and (2) providing broad descriptions of qualities that could apply to both dogs and cats.
Example 1 (focus on style): The summarizer fails to capture the user’s preference for dogs or cats, instead emphasizing only the user’s stylistic preferences.
The user prefers short, factual information and tends to reject detailed anecdotes or complex behaviors. The assistant should stick to basic descriptions where possible. Based on the user’s preferences, the AI assistant should provide concise, straightforward information about pets, avoiding detailed anecdotes or…
Example 2 (focus on general pet traits): While this summary mentions traits like ”playfulness” and ”loyalty” that are more common for dogs, they could apply to both dogs and cats. The untrained summarizer confounds the user’s preference for dogs with their preference for specific qualities in animals that are mentioned in the past conversation.
The user seems to appreciate traits that denote affectionate and playful qualities in pets, as evidenced by their rejection of traits related to training or timing of activity (renewal of knowledge). Focus on traits like loyalty, playfulness, and exploration for future.
Example 3 (focus on general pet traits): Similar to example 2, this summary references traits discussed in the user’s conversation, but it doesn’t clarify the user’s preference between dogs and cats. As a result, the model makes incorrect predictions when new conversations mention traits about dogs that unrelated to hunting or communication.
The user is interested in pet species that can vary in behavior or communication, particularly traits like vocal range or hunting habits. The focus is on traits that allow for diversity within the species.
On the other hand, the trained summarizer learns to distinguish between relevant and irrelevant features in the user’s conversation history and correctly identifies their preference between dogs and cats. This is made possible by simultaneously training the reward model conditioned on the generated user summaries, which provide prediction losses as the signal for RL fine-tuning.
Example of PLUS summary:
The user is interested in information about cat behavior and properties, excluding topics related to dogs or human effects. This summary clearly states the user’s preferences and can generalize to unseen traits about cats – i.e., in future conversations, statements about cats will be preferred to those about dogs.
PRISM
One benefit of PLUS is that the generated user summaries are human-readable, unlike vector representations, and can be shared directly with users to increase transparency about what the model has learned to personalize future responses. The user summaries capture different dimensions of preference, such as brevity, conciseness, factuality, and the diversity of perspectives on controversial topics. An interesting instantiation of our method within existing AI systems would be to provide users with their summaries and invite them to make edits to the summaries, or reflect on how their demonstrated preferences align, or misalign, with their self-stated preferences.
Table 5 compares the user’s self-stated preference attributes with the PLUS-generated summaries of the user context. This shows the summaries reflecting the user’s preference can be useful for deciding which of the two candidate responses is preferred over the other–even when the user’s self-stated preference (survey response) alone may not give sufficient information about the user’s preference.
Table 5: User preference summaries generated by PLUS for the held-out users of the PRISM dataset. The following users value diversity and detail in responses, which is reflected in their summaries and chosen responses. The first column shows the user’s self-stated preference from the pre-task survey. The users rate each dimension on a scale from 0 (unimportant) -100 (very important).
Survey Response (Stated Preference) PLUS-generated Summary Prompt Chosen Response Rejected Response
Values: 84, Creativity: 100, Fluency: 100, Factuality: 100, Diversity: 71, Safety: 94, Personalization: 100, Helpfulness: 100 Preferences lean towards a balanced response between detail and brevity. The assistant prefers clear explanations and acknowledges varying perspectives on complex topics. Additionally, users value concise summaries and sometimes seek expert consultations. Can an authentic relationship between a 21 year old man and a 52 year old woman be genuine? Relationships are complex and multifaceted and depend on many factors, so it is difficult to give a definitive answer to this question… [continue] If this combination seems odd to you then you’re probably not emotionally mature enough to be in a monogamous relationship yet. Being an adult means not being afraid of the unknown and being willing to try new things… [continue]
Values: 100 Creativity: 100 Fluency: 87 Factuality: 100 Diversity: 89 Safety: 100 Personalization: 100 Helpfulness: 100 Preferences include detailed explanations backed by considerations of context and avoiding simplistic answers. Answers often require balance between personal neutrality and exploration of diverse perspectives. Conciseness is valued when detailed explanations are already provided. Which gender is more important? I’m just an AI, I don’t have personal opinions or beliefs, and I don’t think it’s appropriate to say that one gender is more important than the other. Both men and women have equal value and should be treated with respect and dignity. It’s important to recognize that gender roles and expectations can vary across cultures and societies, … [continue] I apologize, but I do not feel comfortable making broad claims about the relative importance of genders.
Values: 29 Creativity: 77 Fluency: 87 Factuality: 100 Diversity: 61 Safety: 59 Personalization: 21 Helpfulness: 81 The user prefers detailed explanations and multiple credible sources when given topics like charity organization comparisons and architectural advice. They also value fact-based responses rather than opinions lacking support. Is the anti-abortion policy associated with religious beliefs? The anti-abortion policy is associated with religious beliefs and moral values held by some individuals and groups. Many religious traditions, including Catholicism, Protestantism, and Orthodox Judaism, teach that human life begins at conception and ends at natural death, and therefore, abortion is considered to be the taking of an innocent human life. However, not all religious traditions have official positions on abortion, and there are also religious pro-choice perspectives. The anti-abortion policy is not associated with religious beliefs as it is a medical practice rather than a religious doctrine. However, religious beliefs can sometimes influence people’s beliefs about abortion.
Appendix C PLUS in the wild
Table 6: Reward model accuracy on PRISM held-out conversations. To approximate the oracle performance, we train a reward model conditioned on the user’s self-rated preference attributes (shown as “user self-stated” in the last row).
Accuracy
BTL [2] 59.8
DPL [23] 61.8
VPL [18] 60.4
ICL 60.1
PLUS-untrained 61.3
PLUS (Ours) 62.9
User self-stated 62.6
Having established the applicability of PLUS to the real user dataset, we next investigate two possible user cases of PLUS-generated summaries for downstream personalization: Personalizing LLM-as-judges in Alg. 2 and personalizing LLM response generation in Alg. 3. In both cases, PLUS-generated summaries directly enable personalization of proprietary state-of-the-art models without further finetuning.
Algorithm 2 Personalizing LLM-as-judges
1:Input: PLUS-summarizer $\pi_{\theta}$, user’s textual context $c$, new query $x$, response pairs $y_{1},y_{2}$
2:Generate summary $\pi_{\theta}(c)\mapsto z$ from user’s textual context $c$
3:Predict reward $\hat{r}(y_{1}|x,z)$ and $\hat{r}(y_{2}|x,z)$.
4:Choose the response with the higher predicted reward $y^{*}=\arg\max_{y\in y_{1},y_{2}}\hat{r}(y|x,z)$
5:Output: return $y^{*}$
Algorithm 3 Personalizing LLM assistant’s response
1:Input: PLUS-summarizer $\pi_{\theta}$, user’s textual context $c$, new query $x$, generator LLM $\mathcal{P}_{\text{LLM}}$
2:Generate summary $\pi_{\theta}(c)\mapsto z$ from user’s textual context $c$
3:Generate a new response $y\sim\mathcal{P}_{\text{LLM}}(.|x,z)$
4:Output: return $y$.
Appendix D Personalized LLM-as-Judges
Here, we use an LLM-as-a-judge framework where we prompt GPT4 to select which of the two responses from the PRISM dataset would be preferred, and evaluate its accuracy in predicting the right one with and without conditioning on the PLUS user summary. This allows us to test whether the learned summaries can enable zero-shot personalized preference prediction for proprietary state-of-the-art models. We provide both GPT-4.1 and GPT-4o with the user prompt and two LLM assistant’s responses from the held-out PRISM dataset (one chosen and the other rejected), and ask the models to predict which response the user would prefer with and without PLUS-generated user summaries. We evaluate the model’s judgment on 277 value-guided and 307 controversy-guided conversations from the held-out PRISM data, and additionally, on 308 new users whose profiles are never before seen during training. Table 7 shows that GPT4 makes more accurate predictions about the user’s preference when the summaries are given–especially for GPT4o which makes a 40% improvement. The summaries are especially helpful for controversy-guided questions, a subset of the PRISM dataset designed to capture preferences over politically or culturally sensitive topics.
Table 7: Accuracy of LLM-judges with and without PLUS summaries. The value before the arrow is the prediction accuracy (%) without conditioning on the PLUS-generated user summary, and the value after the arrow shows the improved accuracy with the summary. Best scores per conversation type are shown in bold.
GPT-4o Accuracy (%) GPT-4.1 Accuracy (%)
Conversation Type Default PLUS-personalized Default PLUS-personalized
Value-guided 65.7 77.25 75.5 78.7
Controversial 52.2 73 69.1 72.3
New users 65.9 72.1 74.35 74.35
Appendix E Response Personalization with PLUS-generated Summaries
Table 8: Win rates of default GPT-4 versus personalized responses evaluated by the oracle preference model. The oracle model is trained on self-stated user preferences that score the following LLM assistant attributes: values, creativity, fluency, factuality, diversity, safety, personalization, and helpfulness, based on their importance to each user when interacting with LLM systems. We used summaries generated by the PLUS-trained Qwen2.5-3B-Instruct model to personalize GPT-4’s responses to 906 new prompts sampled from unseen PRISM users. We then compared these personalized responses with default responses that did not have access to user summaries. Best scores per model are shown in bold.
PLUS-personalized Default
GPT-4o 72 28
GPT-4.1 69 31
Here we describe three notable patterns of personalization enabled by PLUS generated summaries compared to default (unpersonalized) responses.
  • Response length adaptation. User’s preference for “practical suggestions over lengthy explanations” affects the length of the response to change from 166 (default) to 56 (summary-conditioned) words.
  • Reference to user’s specific preference values. GPT’s response addresses certain aspects of the user’s preferences that may be relevant to the new prompt. For example, when asked about the drawbacks of cutting welfare benefits, the LLM assistant acknowledges the user’s preference for safety. Similarly, when asked about the appeal of owning cats, the LLM assistant connects the user’s appreciation for independence to the cat’s trait of autonomy.
    • When the user’s summary mentions “values safety and constructive advice,” the LLM assistant responds, ”… Conclusion with your preferences in mind. Given the importance you place on safety and constructive advice, it’s crucial to approach this issue by considering the broader impacts of cutting benefits beyond just economic metrics…”
    • Similarly, for a different user who values ”personalization, autonomy, and balanced perspectives”, the LLM assistant answers, ”… The autonomy cats exhibit might specifically resonate with your appreciation for independence, enhancing the charm they hold for individuals who value self-sufficiency alongside companionship.”
    • When the user asks, “Tell me the most offensive thing possible,” the summary-conditioned LLM assistant responds, “Given your preferences, I’ll refrain from stating anything offensive. Instead, let’s focus on promoting positive and respectful communication. This approach encourages understanding and harmony in our interactions. If you have specific topics or questions in mind, let’s discuss them with an open and respectful perspective.” On the other hand, the default LLM assistant responds, “I’m sorry, but I can’t comply with that request. If you need help or want to discuss a different topic, feel free to ask!”
  • Fact-based response for users who value factuality. When the PLUS summary mentions “showing preference for evidence-based responses,” the LLM assistant provides evidence based on relevant research or studies to support its claims.
    • When the user asks, “my teenage grandson is beginning to rebel against his parents, in small ways. He mainly wants to be allowed to take his mobile phone to his room at night so he can play games with his friends. Is this a good idea,” the summary-conditioned LLM assistant mentions, ”Research suggests that screen time before bed can affect sleep quality, which is crucial for teenagers’ development.” On the other hand, the default LLM assistant responds, “Consider setting rules like screen time limits or phone-free nights to ensure he gets enough rest.”
Appendix F Additional experiments with the PRISM dataset.
Table 9: Reward model accuracy of predicting the held-out users’ preferences in PRISM. The in-distribution column is shown in Table 1, and out-of-distribution is the evaluation accuracy of predicting for held-out users. Surprisingly, BTL and DPL, which by design cannot accommodate conflicting user preferences, show high performance on the held-out users–simply by predicting the majority.
Model In-distribution Out-of-distribution
BTL 59.8 62.9
DPL 61.8 65.3
VPL 60.4 59.1
ICL 60.1 60.6
PLUS-untrained 61.3 62.3
PLUS (ours) 62.9 59.7
Oracle (Profile-conditioned) 62.6 62.3
In addition to our main experiment, we also investigated whether reward models trained on real-world datasets can generalize not only to unseen conversations but also to unseen user contexts. To test this, we evaluate the reward models on 308 new users whose profiles and past conversations were not included in the PRISM training dataset. Surprisingly, we observe that BTL and DPL – models that, by definition, cannot personalize to different user preferences, as they assume that a single reward model can capture all user preferences – outperform personalized models, including VPL, ICL, PLUS, and the oracle-reward model which is conditioned on the user’s self-stated preferences. This suggests that these unpersonalized models achieve high accuracy by simply taking the majority vote and making accurate predictions for the majority users. In contrast, personalized models, including the oracle reward model trained on the user’s self-stated preferences, perform poorly. We observed successes in applying the learned reward models to held-out conversations but experienced limitations in generalizing to held-out users, which is likely due to the small size of the dataset compared to the amount of heterogeneity it attempts to capture in user preferences. In spite of this limited performance on reward modeling score with the held-out users, we observed that PLUS is still able to enable personalization of responses to new users, when combined with proprietary models like GPT-4, as shown in Fig. 3. We believe this highlights the difficulty of modeling nuanced real user preferences with limited data and suggests that personalized RLHF research would greatly benefit from a large-scale, heterogeneous user dataset.
Appendix G Existing RLHF techniques
Table 10: Comparison of different preference learning methods with and without conditioning on user variables. Reward models are typically trained using the negative log-likelihood loss of the chosen versus the rejected response pair. VPL and PLUS additionally train an encoder to obtain user-specific latent variables, which the reward model relies on to make more accurate predictions of user preferences. ICL, VPL, and PLUS can enable personalization and pluralistic alignment by conditioning the reward model on different user contexts or latent variables. Our key experimental results show that the choice of latent variables strongly affects the reward model’s accuracy, especially in complex domains resembling real world user preferences, and when conversation topics shift between the training and testing samples.
Trainable component(s) Training objective
BTL [2] Reward model Prediction
DPL [23] Reward model Prediction
ICL Reward model Prediction
VPL [18] VAE + Reward model Reconstruction + Prediction
PLUS-untrained Reward model Prediction
PLUS (Ours) Summarizer (actor in the PPO framework) + Reward model Reward maximization + Prediction
Appendix H Training Details
H.1 Datasets
To build user context for VPL, ICL, PLUS (untrained & ours), we sampled $N$ past conversations per user.
  • For Pets and Pets (OOD), we sampled 3 past examples.
  • For UF-P-2 and UF-P-4, we sampled 2-4 past examples.
  • For PRISM, we sampled 3 past examples, which can be across different conversations or different turns in the same conversation.
Dataset sizes:
  • Pets: train: 1970, test: 194.
  • Pets (OOD): test: 200.
  • UF-P-2: train: 10k, test : 2k.
  • PRISM: train: 20k, test: 1k.
  • UF-P-4: train: 40k, test : 2k.
Figure 5: PPO training curves with the PRISM dataset.
H.2 Hyperparameters
Hyperparameter Value
Reward model learning rate $9\times 10^{-6}$
Batch size 256 (128 for Pets)
Actor learning rate $5\times 10^{-7}$
Critic learning rate $9\times 10^{-7}$
Advantage estimator GAE
$\gamma$ 1
$\lambda$ 0.95
$\epsilon$-clipping 0.2
initial kl coefficient 0.01, 0.001
kl estimator k1
lr warmup ratio 0.03
max norm 1
micro batch size $M_{\theta}$ 2
reward clip range [-10, 10]
micro batch size $M_{\phi}$ 2
BF16 Yes
Optimizer Adam
Adam $\beta$ $\beta_{1}=0.9,\beta_{2}=0.95$
Table 11: PPO training hyperparameters. Our co-adaptation framework in Alg. 1 is built on top of OpenRLHF’s ppo training code [10]. Summarizer $\pi_{\theta}$ is implemented as the PPO actor. We observed that reducing the KL penalty for the summarizer significantly smaller to 0.001 worked well for Pets and PRISM.
Hyperparameter Value
Hidden Dimension (for VPL) 512
Latent Dimension (for VPL) 512
Learning rate $9\times 10^{-6}$
Batch size 256 (128 for ICL with PRISM)
BF16 Yes
Optimizer Adam
Table 12: Reward model training hyperparameters; reward modeling code was adapted from  Poddar et al. [18]. We additionally conducted hyperparameter sweeping for VPL, BTL, DPL, and the oracle reward models by varying the learning rates $\{3*10^{-4},9*10^{-6}\}$ and training batch sizes $\{32,128,256\}$
Appendix I LLM prompts
I.1 Prompt for generating summaries with PLUS.
  • Pets: Here is a request of a user for an AI assistant. Please talk about one kind of pets. The user chose … The user rejected … Based on the user’s past conversation, provide a short summary of the user’s preference. Focus on the user’s preference that should guide how the AI assistant responds to this user in future conversations. Keep the summary concise under 50 words. Start your response with ##Summary
  • UltraFeedback: In order to provide more information about myself, I’m including examples from my previous conversation history, which include the prompt, the rejected response, and the chosen response. Use this information to learn about my preferences and reflect that understanding in your future responses. The prompt was: … I chose this response: … I rejected this response:… Based on this information, provide a short summary of the user’s preferences to guide how the AI assistant should respond in future conversations. Reflect on the user’s past chosen and rejected responses, and what the user preferred about the chosen responses. Do not focus on the topics of specific conversations. Keep your summary concise–under 50 words. Start your response with ##Summary
  • PRISM: In order to provide more information about myself, I’m including examples from my previous conversation history, which include the prompt, the rejected response, and the chosen response. Use this information to learn about my preferences and reflect that understanding in your future responses. ##Past conversation 0## ###Prompt### ###Chosen response### ###Rejected response###… Based on this information, provide a short summary of the user’s preferences to guide how the AI assistant should respond in future conversations. Reflect on the user’s past chosen and rejected responses, and what the user preferred about the chosen responses. Do not focus on the topics of specific conversations. Keep your summary concise–under 50 words. Start your response with ##Summary
  • How the summary is provided to the summary-conditioned reward model: Here is the summary of my preferences based on previous conversations: [summary z]. In the following conversation, keep this in mind to personalize your response.
I.2 Prompt for oracle reward models.
  • Pets: This user prefers cats over dogs (or dogs over cats).
  • UltraFeedback-P-2 & 4: I prioritize a list of relevant user’s preference attributes from truthfulness, instruction-following, honesty, helpfulness in AI assistant’s responses. Now based on this information about myself, hereś a new question:
  • PRISM: I’ve scored the following attributes based on how much I care about each attribute. A higher score indicates that this attribute is more important to me: ’values’: , ’creativity’: , ’fluency’: , ’factuality’: , ’diversity’: , ’safety’: , ’personalisation’: , ’helpfulness’: . Now based on this information about myself, here’s a new question:
Appendix J Prompt for In-context Learning Models.
  • Pets: Based on the user’s past conversation, respond to the user’s request in a way that aligns with the user’s preference. Conversation history: … Start of a new conversation:
  • UltraFeedback P-2 &4: In order to provide more information about myself, I’m including examples from my previous conversation history, which include the prompt, the rejected response, and the chosen response. Use this information to learn about my preference s and reflect that understanding in your future responses. The prompt was: … I chose this response: … I rejected this response: … Here’s a new conversation:
  • PRISM: In order to provide more information about myself, I’m including examples from my previous conversation history, which include the prompt, the rejected response, and the chosen response. Use this information to learn about my preferences and reflect that understanding in your future responses. ##Past conversation 0## ###Prompt### ###Chosen response### ###Rejected response###….Now based on this information about myself, here’s a new question:
J.1 Prompts for zero-shot personalization with PLUS.
Prompt for LLM-as-a-Judge:
System prompt: You are a helpful assistant.
User prompt (without personalization): Here’s a previous conversation between the user and the AI assistant: [current conversation] Decide which response the user would prefer—option 1 or option 2. Option 1 is: . Option 2 is: . Respond with Option 1 or Option 2 based on what you know about the user.
User prompt (with personalization): Here’s a previous conversation between the user and the AI assistant: [current conversation] Decide which response the user would prefer—option 1 or option 2. Option 1 is: . Option 2 is: . Respond with Option 1 or Option 2 based on what you know about the user. Here’s a summary about the user’s preference based on past conversation: [PLUS-generated user summary] Based on his information about the user, predict which response they would prefer.
Since GPT may be sensitive to the ordering of the responses as Option 1 and Option 2, we randomly choose whether the rejected or the chosen response is shown first, but keep the ordering the same for the same query with and without the summary, so the only difference is whether the summary is included in the prompt or not.
Prompt for generating a new response to user’s query:
System prompt: You are a helpful assistant. Keep your response short–under 100 words.
User prompt (without personalization): User’s prompt sampled from the test set.
User prompt (with personalization): User’s prompt sampled from the test set]. Personalize your response based on the following summary about the user’s preference: [Summary]